The sum of the ages of a father and his son is 40 years.If they both live until the son becomes as old as the father is now, the sum of their ages will be 96 years . Find their present ages.
Answer by ankor@dixie-net.com(22740)

The sum of the ages of a father and his son is 40 years.If they both live until
the son becomes as old as the father is now, the sum of their ages will be 96 years .
Find their present ages.
:
Let f = father's age
Let s = son's age
:
Write an equation for each statement.
"The sum of the ages of a father and his son is 40 years.'
f + s = 40
s = (40-f)
:
If they both live until the son becomes as old as the father is now, the sum of their ages will be 96 years. "
f + f + (f-s) = 96
2f + (f-s) = 96
3f - s = 96
3f -(40-f) = 96
3f - 40 + f = 96
4f = 96 + 40
4f = 136
f = 136/4
f = 34 is father's age
then
40 - 34 = 6 yrs is son's age
:
:
Check this: the dif in their ages 34-6 = 28;
34 + 28 + 6 + 28 = 96